OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 12 x – y = 3

OpenStudy (anonymous):

To check your answer visit http://www.wolframalpha.com/input/?i=solve+%284x+%2B+2y+%3D+12%2C+++x+%E2%80%93+y+%3D+3++%29

OpenStudy (anonymous):

Multiply the second equation by 2 and add them\[ 6x = 18\\ x=3 \] Using the second equation, you find y=0

OpenStudy (anonymous):

(4x+2y=12)x1 -----1 (x-y=3)x4 ------2 4x+2y=12 -4x+4y=-12 +6y=0 y=0 4x+2(0)=12 4x+0=12 4x=12-0/4 x=3 Check x and y in equation 2 x-y=3 3-0=3

OpenStudy (anonymous):

Thank you.